OpenStudy (anonymous):

Please Solve: (-2+8i)/(-4+4i)

OpenStudy (anonymous):

expand? FOIL just like you would if they were real numbers, remembering that i^2=-1 8-8i-32i+32i^2=8-40i-32=-24-40i

OpenStudy (anonymous):

But dividing wouldnt it be less in fraction?

OpenStudy (anonymous):

oh i didnt see the division sign

OpenStudy (anonymous):

Then it would be... 1+4i/2 ???

sam (.sam.):

\[\frac{-2+8i}{-4+4i}*\frac{-4-4i}{-4-4i}\]

OpenStudy (anonymous):

multiply top and bottom by \[-4-4i\] because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

what .sam. said denominator will be \[4^2+4^2=32\] all the work is multiplying in the numerator

OpenStudy (anonymous):

but the result would be 1+4i/2 ???